For a normal unaided eye the least converging power of the eye lens behind the cornea is 20D and cornea itself has some converging power.The distance between the retina and the cornea - eye lens can be approximated to 5/3 cm. Converging power of the cornea is given by?

Grade:10

1 Answers

Arun
6 years ago

Here,

P1 = P (let) (converging power)

P2 = 20 D

Total Power of the combination  P = P1 + P­2

  = 20+ P

  

Focal length of the combination = 1/(20+P) × 100 cm

  = 100/(20+P)

Now

For minimum converging state of the eye lens object is at infinity

u = -∞

v = 5/3

f = 100/(20+P)

From Lens equation

1/f = 1/v -1/u

=> (20+P)/100 = 3/5 – 1/(-∞)

=> P = 40 D

Hope it helps

 

Regards

Arun(askIITians forum expert)